SubjectCover

Question 13 of 19 in Handling Data 5

Handling Data 5

2 marks · Normal

A fair die is thrown twice. What is the probability of getting at least one six in the two throws?

Work this out on paper, then compare with the answer and worked solution below. With a free account you can submit written answers for marking.

Show answer

11/36 (±0.01)

Also accept: 0.306 (3 sf)

Common mistakes: Forgetting to use 1 – P(no six); Incorrectly using addition of single probability

Show mark scheme and worked solution

Marking points

M1 · 1 mark · Finds probability of no six: 5/6 × 5/6 = 25/36, then complements.

A1 · 1 mark · Gives final answer 11/36 or correct decimal equivalent.

Worked solution

Probability of no six on each throw: 5/6

Probability of at least one six = 1 − 25/36 = 11/36

Report a problem with this question
What's wrong?